OpenStudy (anonymous):

Write the expression in standard form. 7/ 3-15i A. 7/78 - 35/78i B. -7/78 - 35/78i C. 7/78 + 35/78i D. -7/78 + 35/78i

OpenStudy (anonymous):

multiply numerator and denominator by the conjugate of denominator (3+15i). separate real and imaginary parts

OpenStudy (anonymous):

so would the answer be C?

OpenStudy (anonymous):

i'm confused :(

OpenStudy (anonymous):

u r right.. dnt confuse

OpenStudy (anonymous):

so it is C.. ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ohh ok thankyou!! by any chance would u no the answer to this.. Find the product of the complex number and its conjugate. 1 + 3i A. 1 + 9i B. 10 C. -8 D. 1 - 9i i kept getting 1-9i .. but im not sure?

OpenStudy (anonymous):

(1+3i)x(1-3i) is of the form (a+b)(a-b)=a^2-b^2 here a=1 and b=3i so a^2 - b^2=_________

OpenStudy (anonymous):

= 10?

OpenStudy (anonymous):

yep